OpenStudy (anonymous):

pleaseeeeeeeeee helppppp A $15,000, 11 percent, 120-day note, dated Sept. 3, is discounted on Nov. 11. Assuming a bank discount rate of 9 percent, the proceeds would be____________. 1. $15,550.00 2. $15,351.74 3. $15,531.74 4. $15,135.47

OpenStudy (anonymous):

we use formula ( Total - initial)/ ( initial ) = discount rate

OpenStudy (anonymous):

I keep getting $15,508.92. But I'm guessing at a lot of the meanings.

OpenStudy (anonymous):

first find the maturity value of the note, 15000+[15000*0.11*120/360]=15,550 then find the term of discount,120-[27+31+11]=51days then multiply the maturity value by discount rate and the term of discount, 15,550*0.09*51/360=198.26 then get the difference between the maturity value and the discount,15550-198.26=15,351,74 so the proceeds is 15,351.74 which is #2